Service has been fully restored to the Summit area. If you are still experiencing issues, please reset your connection by unplugging and then plugging back in your personal router, Plume POD(s), or VNET-provided Calix router. For additional assistance, call us at 814-636-1500.
For all other areas, including Harborcreek, our team is actively working to resolve the situation and expects a resolution shortly.
Posted Mar 21, 2025 - 16:03 EDT
Update
Service restoration efforts are ongoing. Technicians are currently working in Harborcreek to restore service, while our engineers are in the final stages of remediation for the Summit area.
Posted Mar 21, 2025 - 13:41 EDT
Update
We are actively working on a remediation for the outages currently affecting certain areas, particularly in the Summit and Harborcreek areas. Engineers and technicians are working diligently to restore service as quickly as possible.
Posted Mar 21, 2025 - 11:09 EDT
Update
Crews are provisioning a replacement card and will be dispatched to our Summit location to install, which will stabilize services in the area
Posted Mar 21, 2025 - 06:55 EDT
Update
We are investigating reports of users in the Summit Township area being disconnected. Network team is investigating and dispatching crews to the area.
Posted Mar 21, 2025 - 06:14 EDT
Identified
Our engineers have identified the issue and are working on a resolution. The affected locations are now coming back online. If you are not connecting to the internet, please reset power to your router (if you have provided your own) or your Plume POD / Calix router plugged into your modem.
Posted Mar 21, 2025 - 04:47 EDT
Update
After consulting with routing vendors, we will be performing emergency maintenance on network between 1:30am - 2:30am. Connectivity may be disrupted during this time. Please do not reboot any of your equipment as it will recover after the maintenance. Further updates will be provided after the maintenance.
Posted Mar 21, 2025 - 01:24 EDT
Update
Engineering teams believe to have found the source of the issue and are working with vendors to mitigate is in progress.
Posted Mar 21, 2025 - 00:29 EDT
Update
Engineering teams are continuing to assess the connectivity issue affecting a segment of fiber customers with our software vendors. More updates will be provided as engineering teams are provided more information.
Posted Mar 20, 2025 - 22:24 EDT
Investigating
Please note that your VNET Fiber service may be affected by an outage in your area. Our network team is currently working to resolve the issue. You do not need to reboot any equipment and service will come back online automatically. We apologize for the inconvenience.